# High Accuracy, Less Talk ($\mathrm{HALT}$): Reliable LLMs through Capability-Aligned Finetuning

### v1

@@ Abstract
Large Language Models (LLMs) currently respond to every prompt. However, they can produce incorrect answers when they lack knowledge or capability – a problem known as hallucination. We instead propose post-training an LLM to generate content only when confident in its correctness and to otherwise (partially) abstain. Specifically, our method, $\mathrm{HALT}$, produces capability-aligned post-training data that encodes what the model can and cannot reliably generate. We generate this data by splitting responses of the pretrained LLM into factual fragments (atomic statements or reasoning steps), and use ground truth information to identify incorrect fragments. We achieve capability-aligned finetuning responses by either removing incorrect fragments or replacing them with “Unsure from Here” – according to a tunable threshold that allows practitioners to trade off response completeness and mean correctness of the response’s fragments. We finetune four open-source models for biography writing, mathematics, coding, and medicine with $\mathrm{HALT}$ for three different trade-off thresholds. $\mathrm{HALT}$ effectively trades off response completeness for correctness, increasing the mean correctness of response fragments by 15% on average, while resulting in a 4% improvement in the F1 score (mean of completeness and correctness of the response) compared to the relevant baselines. By tuning $\mathrm{HALT}$ for highest correctness, we train a single reliable Llama3-70B model with correctness increased from 51% to 87% across all four domains while maintaining 53% of the response completeness achieved with standard finetuning.
[…]
Tim Franzmeyer , Madian Khabsa
@@ 2.1 Creating a capability-aligned finetuning dataset with $\mathrm{HALT}$
For each prompt $x_{j}$, we first generate a response that aligns with the format of the ground truth responses in the finetuning dataset $D$ through few-shot prompting of the pretrained model. We sample a set of four prompt-response pairs $\{(x_{k},y_{k})\}_{k=1}^{4}$ uniformly at random from $D\setminus\{(x_{j},y_{j})\}$. The context $C_{j}=\left\{(x_{k},y_{k})\right\}_{k=1}^{4}$ is concatenated with the target prompt $x_{j}$ using a question-answer format. The pretrained model $\mathcal{M}$ is then prompted to create a preliminary response $y_{j}^{\text{pt}}$ for the prompt $x_{j}$, i.e. $y_{j}^{\text{pt}}=\mathcal{M}(\text{concat}(C_{j},x_{j}))$, where concat denotes the concatenation of the context $C_{j}$ and the target prompt $x_{j}$ into a single input sequence.
@@
2.2 Trading off between Response Correctness and Completeness
3 Experimental Validation
Overview.
3.1 Implementation and Validation of Relevant Components
Prior Work and Baselines.
Datasets.
Splitting Responses into Fragments.
Evaluator Implementation.
For Wikibios, we prompt Llama3-405B with the fragment to assess, as well as with the entire Wikipedia article of the relevant person, and ask it to assess whether the given fragment is correct or incorrect in the given context (see example prompt in App. B.1). Hence, the evaluator is called multiple times to assess a single response. For MATH, MedExQA, and Wikibios, we prompt Llama3-405B with the numbered fragments that make up the response and the ground-truth step-by-step solution, and prompt it to identify the first incorrect fragment.
F1 Score to Measure the Mean of Completeness and Correctness.
@@ 3.1 Implementation and Validation of Relevant Components
We evaluate the performance of our evaluator in assessing the correctness of individual fragments of responses of all four datasets. For each dataset, we randomly sampled 300 responses generated by Llama3-70B and manually annotated the correctness of all fragments. We observe in Table 1 that the Llama3-405B evaluator’s error ranges from 0.27 incorrectly assessed atomic statements per Wikipedia response, composed of 9.58 atomic statements on average, to a discrepancy of 1.14 lines on Coding, with an average number of 7.85 lines of code per response. In MATH, the evaluator misjudges the location of the first incorrect fragment by only 0.6 reasoning steps, at an average response length of 5.54. We conclude that Llama3-405B – prompted with relevant information – is well suited to assess the correctness of individual logical fragments.
@@ 3.3 Training a Multi-Domain LLM Achieving 87% Correctness
Instead of only being provided with response fragments that the model is certain about, users might desire to observe the model’s full response to a prompt while being informed which parts of the response might be incorrect. This is in contrast to responses achieved with standard finetuning, where the user is not informed which parts (fragments) the model is uncertain about. For example, this would enable users to edit generated Code which the LLM is uncertain about or to verify correctness of specific statements using external sources. We evaluate this use case by modifying $\mathrm{HALT}$ to annotate uncertain response fragments with “Uncertain”, instead of omitting them. Specifically, we modify the finetuning dataset for Math, Medical Q., and Coding by marking the first incorrect response fragment and all subsequent fragments as “Uncertain”. Note that these fragments are replaced by “Unsure from here” in the default $\mathrm{HALT}$ response. We do not investigate the open-ended Wikipedia biography writing task as it allows for generating an unbounded number of “Uncertain” statements. Figure 6 in the Appendix shows that training Llama3-70B at $\alpha=60\%$ with Add-On HALT results in significant correctness improvements when fragments marked as “Uncertain” are excluded from the correctness evaluation, achieving improvements similar to those of $\mathrm{HALT}$’s default implementation. When the “Unsure” markers are ignored Add-On HALT achieves response completeness similar to that achieved when finetuning on the Unchanged responses. In conclusion, a slight modification of HALT enables the retention of response completeness while reliably informing the user about likely incorrect fragments.
@@ 4 Related Work
Another line of prior work has found that LLMs are well-calibrated, and found that confidence scores can, for example, be obtained via inspection of logits (Kadavath et al., 2022) or via prompting LLMs to state their confidence (Lin et al., 2022). This provides a basis for recent works that train models to abstain when uncertain (Zhang et al., 2024a; Chen et al., 2023; Yadkori et al., 2024; Wen et al., 2024; Brahman et al., 2024; Feng et al., 2024; Cheng et al., 2024; Tuan et al., 2024). Unlike these works, which train models to either fully abstain or fully answer, which is unsuitable for prompts that require long responses, $\mathrm{HALT}$ trains LLMs to selectively compose responses according to the given capabilities. In further contrast, $\mathrm{HALT}$ enables modification of the desired trade off between completeness and correctness of responses.
@@ 5 Conclusion
We present HALT, a novel finetuning paradigm that trains models to generate responses according to their internal confidence. HALT relies on evaluating the knowledge and capabilities of a given pretrained model with respect to specific prompts and trains the pretrained model to omit response fragments it is uncertain about. HALT training yields models that are up to 37% more accurate than those trained using unchanged ground-truth answers while maintaining high response completeness.
Limitations and Future Work.
Future work could address the following limitations. First, our method assumes either independent factual fragments or causal dependence. This could be addressed by instead establishing dependency graphs; for instance, coding responses might be decomposed into smaller blocks whose correctness depends on other blocks. Second, $\mathrm{HALT}$ relies on a binary assessment of whether a fragment is within the model’s capability, but a more fine-grained scale and thresholding could improve performance. Third, we train separate models for different correctness–completeness trade-offs. Training a single model that adapts its responses based on a user-selected parameter would improve flexibility. Fourth, our approach assumes no access to external sources. Instead, LLMs should call tools or acquire external information when uncertain about the remaining response. Fifth, we treat all factual fragments as equally valuable, though weighting them by their impact on the final answer could improve utility. Finally, we focus on text-based models and datasets; extending $\mathrm{HALT}$ to multi-modal domains is a promising direction.
@@ A.2 Additional Implementation Details
We use supervised finetuning with LoRA (Hu et al., 2021) for compute-efficient training, with a rank of 256 and alpha of 512. Training is performed using the AdamW (Kingma, 2014) optimizer with betas set to $(0.9,0.95)$, weight decay of $0.01$, and an initial learning rate of $51e-$ 5. The learning rate schedule follows a warmup fraction of 5%, with a decay to $61e-$ 6 over 5 epochs. The effective batch size is set to 128.
@@ B.1 Example Evaluator Prompts
[…]
@@
Example for $\alpha=0.4$*** Question:****
@@
Example for $\alpha=0.6$*** Question:****
@@
Example for $\alpha=0.8$*** Question:****
@@
Example for $\alpha=0.4$*** Question:****
@@
[…]
Example for $\alpha=0.6$*** Question:****
@@
[…]
Example for $\alpha=0.8$*** Question:****
@@
[…]

### v2

# High Accuracy, Less Talk ($\mathrm{HALT}$): Reliable LLMs via Capability-Aligned Finetuning

@@ Abstract
Large Language Models (LLMs) currently respond to every prompt. However, they can produce incorrect answers when they lack knowledge or capability – a problem known as hallucination. We instead propose post-training an LLM to generate content only when confident in its correctness and to otherwise (partially) abstain. Specifically, our method, $\mathrm{HALT}$, produces capability-aligned post-training data that encodes what the model can and cannot reliably generate. We generate this data by splitting responses of the pretrained LLM into factual fragments (atomic statements or reasoning steps), and use ground truth information to identify incorrect fragments. We achieve capability-aligned finetuning responses by either removing incorrect fragments or replacing them with “Unsure from Here” – according to a tunable threshold that allows practitioners to trade off response completeness and mean correctness of the response’s fragments. We finetune four open-source models for biography writing, mathematics, coding, and medicine with $\mathrm{HALT}$ for three different trade-off thresholds. $\mathrm{HALT}$ effectively trades off response completeness for correctness, increasing the mean correctness of response fragments by 15% on average. By tuning $\mathrm{HALT}$ for highest correctness, we train a single reliable Llama3-70B model with correctness increased from 51% to 87% across all four domains while maintaining 53% of the response completeness achieved with standard finetuning.
@@ 1 Introduction
We remark that no model will ever achieve perfect performance across all possible tasks, requiring capability-aligned responses regardless of model strength (OpenAI, 2025). Additionally, $\mathrm{HALT}$ is not about distilling from a stronger model; rather, it directly addresses the problem of ensuring that an LLM’s outputs are aligned with its own uncertainty.
@@ 2.1 Creating a capability-aligned finetuning dataset with $\mathrm{HALT}$
For each prompt $x_{j}$, we first generate a response that aligns with the format of the ground truth responses in the finetuning dataset $D$ through few-shot prompting of the pretrained model. We sample a set of four prompt-response pairs $\{(x_{k},y_{k})\}_{k=1}^{4}$ uniformly at random from $D\setminus\{(x_{j},y_{j})\}$ as in-context learning examples. The context $C_{j}=\left\{(x_{k},y_{k})\right\}_{k=1}^{4}$ is concatenated with the target prompt $x_{j}$ using a question-answer format. The pretrained model $\mathcal{M}$ is then prompted to create a preliminary response $y_{j}^{\text{pt}}$ for the prompt $x_{j}$, i.e. $y_{j}^{\text{pt}}=\mathcal{M}(\text{concat}(C_{j},x_{j}))$, where concat denotes the concatenation of the context $C_{j}$ and the target prompt $x_{j}$ into a single input sequence.
@@
2.2 Trading off between Response Correctness and Completeness
3 Experimental Validation
Overview.
3.1 Implementation and Validation of Relevant Components
Prior Work and Baselines.
Datasets.
Splitting Responses into Fragments.
Evaluator Implementation.
For Wikibios, we prompt Llama3-405B with the fragment to assess, as well as with the entire Wikipedia article of the relevant person, and ask it to assess whether the given fragment is correct or incorrect in the given context (see example prompt in App. B.1). Hence, the evaluator is called multiple times to assess a single response. For MATH, MedExQA, and APPS, we prompt Llama3-405B with the numbered fragments that make up the response and the ground-truth step-by-step solution, and prompt it to identify the first incorrect fragment.
F1 Score to Measure the Mean of Completeness and Correctness.
@@ 3.1 Implementation and Validation of Relevant Components
We evaluate the performance of our evaluator in assessing the correctness of individual fragments (denoted as $f_{1},\dots,f_{k}$ for a response split into $k$ fragments, as defined in Section 4) of responses of all four datasets. For each dataset, we randomly sampled 300 responses generated by Llama3-70B and manually annotated the correctness of all fragments. We observe in Table 1 that the Llama3-405B evaluator’s error ranges from 0.27 incorrectly assessed atomic statements per Wikipedia response, composed of 9.58 atomic statements on average, to a discrepancy of 1.14 lines on Coding, with an average number of 7.85 lines of code per response. In MATH, the evaluator misjudges the location of the first incorrect fragment by only 0.6 reasoning steps, at an average response length of 5.54. We conclude that Llama3-405B – prompted with relevant information – is well suited to assess the correctness of individual logical fragments.
@@ 3.2 HALT Enables Trading off Completeness and Correctness
HALT Does Not Harm General Instruction-Following Capabilities.
To assess whether HALT maintains coherence and general instruction-following capabilities beyond domain-specific accuracy, we conducted an AlpacaEval-style evaluation (Li et al., 2023) comparing Llama3-70B HALT (trained with $\alpha=0.6$) against the base Llama3-70B model across 578 prompts spanning all four evaluation datasets, using GPT-5 as a judge. HALT achieves near-parity with the base model overall (50.8% win rate), demonstrating that selective abstention does not significantly harm general instruction-following capabilities. Notably, HALT outperforms the base model on wikibios_v2 (66.0% win rate), where factual accuracy is critical—precisely the use case HALT is designed for. On mathematical reasoning and general-domain tasks (MATH, appsintro, medex), HALT shows modest degradation (41.8-45.4% win rates), expected since judges may penalize abstentions when completeness is valued over guaranteed correctness (see Table 5 in the Appendix for full results).
@@ 3.3 Training a Multi-Domain LLM Achieving 87% Correctness
Users might desire to observe the model’s full response while being informed which parts might be incorrect, enabling them to edit uncertain code or verify statements externally. We evaluate this use case by modifying $\mathrm{HALT}$ to annotate uncertain response fragments with “Uncertain”, instead of omitting them. Specifically, we modify the finetuning dataset for Math, Medical Q., and Coding by marking the first incorrect response fragment and all subsequent fragments as “Uncertain”. Note that these fragments are replaced by “Unsure from here” in the default $\mathrm{HALT}$ response. We do not investigate the open-ended Wikipedia biography writing task as it allows for generating an unbounded number of “Uncertain” statements. Figure 6 in the Appendix shows that training Llama3-70B at $\alpha=60\%$ with Add-On HALT results in significant correctness improvements when fragments marked as “Uncertain” are excluded from the correctness evaluation, achieving improvements similar to those of $\mathrm{HALT}$’s default implementation. When the “Unsure” markers are ignored Add-On HALT achieves response completeness similar to that achieved when finetuning on the Unchanged responses. In conclusion, a slight modification of HALT enables the retention of response completeness while reliably informing the user about likely incorrect fragments.
@@ 4 Related Work
Another line of prior work has found that LLMs are well-calibrated, and found that confidence scores can, for example, be obtained via inspection of logits (Kadavath et al., 2022) or via prompting LLMs to state their confidence (Lin et al., 2022). This provides a basis for recent works that train models to abstain when uncertain (Zhang et al., 2024a; Chen et al., 2023; Yadkori et al., 2024; Wen et al., 2024; Brahman et al., 2024; Feng et al., 2024; Cheng et al., 2024; Tuan et al., 2024). Recent work has explored training models to express verbalized uncertainty through confidence scores (Band et al., 2024) and calibrated hedging (Stengel-Eskin et al., 2024). Unlike prior works that either fully abstain, fully answer, or preserve complete responses with uncertainty qualifications, $\mathrm{HALT}$ trains LLMs to selectively compose partial responses according to their capabilities, enabling modification of the desired trade off between completeness and correctness.
@@ 5 Conclusion
We present HALT, a novel finetuning paradigm that trains models to generate responses according to their internal confidence by omitting response fragments they are uncertain about. HALT training yields models that are up to 37% more accurate than those trained using unchanged ground-truth answers while maintaining high response completeness.
Limitations.
Our method assumes independent or causally dependent fragments; dependency graphs could improve complex tasks. HALT uses binary capability assessment; fine-grained scaling may help. We train separate models per trade-off; a single adaptive model would improve flexibility. Extending to multi-modal domains is promising.
@@ A.2 Additional Implementation Details
We use supervised finetuning with LoRA (Hu et al., 2021) for compute-efficient training, with a rank of 256 and alpha of 512. Training is performed using the AdamW (Kingma, 2014) optimizer with betas set to $(0.9,0.95)$, weight decay of $0.01$, and an initial learning rate of 1e51e-5. The learning rate schedule follows a warmup fraction of 5%, with a decay to 1e61e-6 over 5 epochs. The effective batch size is set to 128.
@@
A.3 AlpacaEval General Instruction-Following Results
Dataset HALT Wins Base Wins Ties Win Rate
MATH 49 59 42 45.4%
appsintro 46 64 39 41.8%
medex 54 65 9 45.4%
wikibios_v2 97 50 3 66.0%
Overall 246 238 93 50.8%
Table 5: AlpacaEval-style evaluation comparing HALT ($\alpha=0.6$) against base Llama3-70B across 578 prompts. Win rate excludes ties. GPT-4o-mini served as judge.
@@ B.1 Example Evaluator Prompts
Example prompt to Llama3-405B evaluator to determine correctness of a response composed of causally dependent fragments. ****Math Question:**** A triangle has a base of 8 units and a height of 5 units. If the triangle is part of a rectangle with the same base and height, what fraction of the rectangle’s area does the triangle occupy? ****Ground truth Response for Math Question:**** The area of a triangle is calculated using the formula: $\text{Area}=\frac{1}{2}\times\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=\frac{1}{2}\times 8\times 5=20\ \text{square units}$. The area of the rectangle is calculated using the formula: $\text{Area}=\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=8\times 5=40\ \text{square units}$. The triangle occupies a fraction of the rectangle’s area given by: $\frac{\text{Triangle's area}}{\text{Rectangle's area}}=\frac{20}{40}=\frac{1}{2}$. Thus, the triangle occupies $\frac{1}{2}$ of the rectangle’s area. ****Step-wise Response given by student:**** Line 1: The area of a triangle is calculated using the formula: $\text{Area}=\frac{1}{2}\times\text{base}\times\text{height}$. Line 2: Substituting the given values: $\text{Area}=\frac{1}{2}\times 8\times 5=20\ \text{square units}$. Line 3: The area of the rectangle is $8\times 5=40\ \text{square units}$. Line 4: The fraction of the rectangle’s area that the triangle occupies is 20square units20\ $\text${square units}. ****Instruction:**** You are given a Math question above, as well as a Ground truth Response to it. Determine until which line the step-wise response given by the student is correct. Reply with the number of the last correct line. If the first step is already wrong, reply with zero. Otherwise, reply with the number of the last correct line, i.e., with an integer between 0 and the number of lines in the response. Reply with ’Line X’ only, where X is the number of the last correct line.
@@
Example for $\alpha=0.4$ **** Question:****
@@
Example for $\alpha=0.6$ **** Question:****
@@
Example for $\alpha=0.8$ **** Question:****
@@
Finetuning pair for $\alpha=0.4$ **** Question:****
@@
**** Response:**** The area of a triangle is calculated using the formula: $\text{Area}=\frac{1}{2}\times\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=\frac{1}{2}\times 8\times 5=20\ \text{square units}$. Unsure from here.
Finetuning pair for $\alpha=0.6$ **** Question:****
@@
**** Response:**** The area of a triangle is calculated using the formula: $\text{Area}=\frac{1}{2}\times\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=\frac{1}{2}\times 8\times 5=20\ \text{square units}$. The area of the rectangle is calculated using the formula: $\text{Area}=\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=8\times 5=40\ \text{square units}$. Unsure from here.
Finetuning pair for $\alpha=0.8$ **** Question:****
@@
**** Response:**** The area of a triangle is calculated using the formula: $\text{Area}=\frac{1}{2}\times\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=\frac{1}{2}\times 8\times 5=20\ \text{square units}$. The area of the rectangle is calculated using the formula: $\text{Area}=\text{base}\times\text{height}$. Substituting the given values: $\text{Area}=8\times 5=40\ \text{square units}$. The triangle occupies a fraction of the rectangle’s area given by: $\frac{\text{Triangle's area}}{\text{Rectangle's area}}=\frac{20}{40}=\frac{1}{2}$. Thus, the triangle occupies $\frac{1}{2}$ of the rectangle’s area.